a rectangle has a perimeter of 12 m. If each side is a whole number of meters, what are the possible dimensions for the length and width? List them and explain my answer.
Answer by checkley77(12569)

2L+2W=12
2L**2W**=12
2*1+2*5=12
2*2+2*4=12
2*3+2*3=12